Solve the inequality
(-6x+2)/(3x^2+7) >0

Answer by RAY100(1637)

(-6x+2) (3x^2+7) > 0
.
lets solve eqn as if it were equal to zero
.
(-6x+2) (3x^2 +7) = 0
.
-6x+2 = 0
.
-6x = -2
.
x= + 1/3,,,,,zero
.
3x^2 +7 =0
.
3x^2 =-7
x^2 = -7/3
x= +/- 1.53 i,,,,imaginary therefore not a Real Number Solution
.
The only Real answer is x= +1/3
.
If we had solved this part of the problem with the inequality problem with the inequality sign, -6x+2 >0, the answer would be
x < 1/3.
.
Lets check by picking test points on either side of + 1/3
.
Smaller than + 1/3, use 0,
.
substituting,,,(-6(0)+2)(3(0)^2 +7) = 2*7=14,,,,or 14 > 0 ,,,,ok
.
Larger than +1/3 ,,,use 1
.
subst,,,,,(-6(1)+2)(3(1)^2 +7) = (-4)(10) = -40,,,,or -40>0,,,,not true
.
Therefore ,,,,,x< 1/3,,,,is valid answer






.